# Modeling of acceleration in heavy-ion collisions: occurrence of temperature below the Unruh temperature

### v1

@@ I Introduction
Since the beginning of the 21st century, the field of research of relativistic heavy ion collisions has undergone rapid development, initiated by the active operation of the Relativistic Heavy Ion Collider (RHIC) and, later, of a related program at the Large Hadron Collider (LHC). The discovery of a new state of nuclear matter [1, 2, 3, 4], known as the quark-gluon plasma (QGP), has prompted a rethinking of the initial assumptions regarding the behavior of high-temperature quantum chromodynamics (QCD) matter. Specifically, the prevailing expectations of a weakly coupled gas-like state [5, 6, 7, 8] have been challenged, suggesting instead a behavior more akin to a strongly-coupled almost ideal fluid [9, 10, 11, 12, 13].
@@
From the point of view of hydrodynamics, the direct analogue of the magnetic field $B_{\mu}=\frac{1}{2}\varepsilon_{\mu\nu\alpha\beta}u^{\nu}F^{\alpha\beta}$ is vorticity $\omega_{\mu}=\frac{1}{4}\varepsilon_{\mu\nu\alpha\beta}u^{\nu}\Omega^{\alpha\beta}$, and the direct analogue of the electric field $E_{\mu}=F_{\mu\nu}u^{\nu}$ is acceleration $a_{\mu}=-\Omega_{\mu\nu}u^{\nu}$ (here $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ is the usual electromagnetic field tensor, and $\Omega_{\mu\nu}=\partial_{\mu}u_{\nu}-\partial_{\nu}u_{\mu}$ is the kinematic vorticity tensor, while $u^{\mu}$ is the four-velocity of the medium), being, respectively, the vector ($E_{\mu}$ and $a_{\mu}$) and pseudovector ($B_{\mu}$ and $\omega_{\mu}$) components of corresponding tensors ($F_{\mu\nu}$ and $\Omega_{\mu\nu}$). And while vorticity and electric fields have been studied in details, there are not many works devoted to modeling acceleration (see, for example [70, 71]).
The present research aims to examine the process of generation and evolution of acceleration in both central and non-central heavy-ion collisions simulated within the PHSD model. The focus is on modeling spatial distributions of acceleration at different time moments of the system evolution, with the intention of evaluating acceleration (and the corresponding Unruh temperature) and comparing it with the temperature from the equation of state (EoS). The existence of extremely high accelerated zones, consisting of hadrons, where the Unruh temperature exceeds that obtained from the EoS, will be demonstrated. Furthemore, it will be shown that the partonic phase is characterized by minimal acceleration, which results in the temperature being below the Unruh temperature  [^1] Let us emphasize that the difference between the temperature and the Unruh temperature is not at all connected with any “violation” of the Unruh effect, as it might naively seem. According to the Unruh effect, the temperature is equal to the Unruh temperature only for the concrete quantum (vacuum) state, as discussed below..
@@ II Calculating acceleration and temperature within PHSD
×$\quad\times\delta\big{(}p_{0}-\sqrt{m_{h}^{2}+\boldsymbol{p\,}^{2}}\big{)}\delta^{(3)}\big{(}\boldsymbol{r}-\boldsymbol{r}_{i_{h}}(t)\big{)},$ (3)
@@
×$\quad\times\delta\big{(}p_{0}-\sqrt{m_{h}^{2}+\boldsymbol{p\,}^{2}}\big{)}\Phi(\boldsymbol{r},\boldsymbol{r}_{i_{h}}(t)),$ (4)
@@
To determine the local temperature, we use the EoS [87] and solve $T(n_{B},\varepsilon)=T$, where $\varepsilon$ is taken from (6) and $n_{B}$ is the baryon density, calculated from the baryon current $J^{\mu}_{B}$ as $n_{B}=u_{\mu}J^{\mu}_{B}$. Details of the algorithm used, the explicit form of the smearing kernel $\Phi$, the spectator separation procedure, the analysis of the resulting medium, and the application to the calculation of the global polarization can be found in a series of papers [56, 88, 89, 90].
@@
The PHSD transport model employs the parallel ensemble method, which involves the simultaneous simulation processing of $N$ collisions. To improve the accuracy of the calculated values, the PHSD model is initialized $M$ times, with the fluidization procedure being performed for each run individually. Consequently, the total number of simulated events is $N\times M$. In this paper, the number of parallel ensembles was set to $N=200$ for the collision energies of $7.7$ and 11.511.511.5\,11.5GeV, and the number of independent model runs is $M\approx 500$. Thus, the total statistics is $(5-10)\times 10^{4}$ events for each impact parameter and collision energy.
@@ III Results
In this sections, we present a detailed analysis of temperature $T$ (obtained from the EoS), the Unruh temperature $T_{\rm U}$ (calculated as (1) and (7)), and the difference $T^{2}-T_{\rm U}^{2}$ for the medium produced in the central ($b=0\,$fm) and off-center ($b=7.5\,$fm) heavy-ion collisions.
@@ III.1 Temperature and acceleration in central and off-center collisions: overall analysis
Figure 1: Profiles of temperature $T$ (first column), characteristic Unruh temperature $T_{\rm U}$ (second column) in the $z=0$ plane, calculated at different times after the collision for Au + Au collisions at $\sqrt{s_{NN}}=7.7\,$GeV. The profiles of $T$ and $T_{\rm U}$ shown in (a) are calculated for a central collision ($b=0\,$fm); in (b) for an off-center collision ($b=7.5\,$fm). The dotted curves show two contours in the $xy$ plane, limiting the medium in energy density to the value $\varepsilon=500\,$MeV/fm3 (inner contour) and $\varepsilon=50\,$MeV/fm3 (outer contour).
@@
The time $t-t_{coll}=0\,$fm$/c$ corresponds to the moment of contact of the nuclei. At this moment, a hot medium begins to form in the center of the system, which then expands. The volume occupied by the hot substance changes non-trivially [89], and the temperature of the resulting medium reaches its highest value at the moment of maximum overlap of the nuclei ($t-t_{coll}\approx 2.5\,$fm$/c$). The temperature profiles $T$ in Figure 1 show that the hottest medium in the central cut is formed in the time interval from 2 to 3 fm$/c$ after the moment of contact of the nuclei for the collision under consideration. After 3 fm$/c$, the resulting medium cools down due to hadronization and thermalization processes [14, 76, 77, 78]. As can be seen from the temperature profiles $T$, the hot medium in the plane $z=0\,$fm is located inside the contour that limits the medium in energy density to the value $\varepsilon=50\,$MeV/fm3. In what follows, by a fireball we mean a medium limited by this energy density ($\varepsilon\geq 50\,$MeV/fm3). In the PHSD model, the QGP is defined to exist inside a region limited by the energy density to a value of $\varepsilon=500\,$MeV/fm3 (see [84] and ref. therein). In Figure 1 we denoted this region by dashed black and white contour. Note that such an analysis is valid for a short period of time after the collision ($t-t_{coll}=3-4\,$fm$/c$). In the PHSD model, after the maximum overlap of nuclei the fraction of the partonic phase rapidly decreases and becomes insignificant [91]. As can be seen in Figure 1, the temperature $T$ has reaches a maximum in the region of QGP phase and is about 200 MeV for the time of maximal overlap in the central cut.
@@
The characteristic Unruh temperature, as can be seen from the $z=0\,$fm profiles in Fig. 1, has a maximal value at early times after the nuclei touch $t-t_{coll}=0.5\,$fm$/c$. Then the value of the characteristic Unruh temperature rapidly decreases with the evolution and expansion of the resulting medium. The most accelerated medium in the central cut is formed mainly outside the fireball, where the medium consists of hadrons. In the QGP region, $T_{\rm U}$ has a very small value at the initial time after the collision and practically disappears by the time of maximum overlap. Thus, in the central cut the cold regions outside the fireball, consisting of hadrons, are highly accelerated. The received pictures exhibit core-corona structures.
@@
The qualitative picture of the evolution of the temperature $T$ and the characteristic Unruh temperature $T_{\rm U}$ remains the same for both central collisions ($b=0\,$fm, Fig. 1(a)) and off-center ones ($b=7.5\,$fm, Fig. 1(b)). The main difference is that for the non-central collisions the medium volume decreases and the acceleration is slightly higher than for the central ones.
@@ III.2 Comparison: partonic and hadronic phases
In order to better understand the physical picture presented in Figure 1 at a qualitative level, we calculated a time evolution of the temperatures $T$ and $T_{\rm U}$ separately for partonic and hadronic phases, and averaged over the entire volume of the fireball. Such plots for Au + Au collisions at energies $\sqrt{s_{NN}}=4.5-11.5\,$GeV are presented in Figure 2.
@@
Figure 2: Evolution of the temperature $T$ and the characteristic Unruh temperature $T_{\rm U}$, calculated in PHSD for Au + Au collisions at energies $\sqrt{s_{NN}}=4.5,7.7$ and 11.511.511.5\,11.5GeV. For the collision energy $\sqrt{s_{NN}}=7.7\,$GeV, two impact parameters are considered: $b=0\,$fm, which corresponds to a central collision, and $b=7.5\,$fm, which corresponds to an off-center collision of nuclei. The results are presented for a full fireball $\varepsilon>50\,$MeV/fm3 (first panel); for partonic phase only (second panel) and for hadronic phase only (third panel).
@@
Let us first discuss the results obtained for the collision energies $\sqrt{s_{NN}}=7.7$ and 11.511.511.5\,11.5GeV for the fireball (Fig. 2 (first panel)). After the moment of contact of the nuclei, when the fireball is just beginning to form, the characteristic Unruh temperature has a maximum value. Then the value of $T_{\rm U}$ decreases rapidly, which is associated with the rapid expansion of the fireball. The temperature $T$ of the fireball has a maximum value at the moment of maximum overlap of the nuclei, which corresponds to $t-t_{coll}\approx 2.5\,$fm$/c$. Then the medium begins to cool, which is associated with the processes of hadronization and thermalization. Note that for the collision energies $\sqrt{s_{NN}}=7.7$ and 11.511.511.5\,11.5GeV, the obtained maximum value of the temperature $T$ of the resulting medium at the moment of maximum overlap is $T\approx 160-170\,$MeV, which is close to the temperature of the confinement-deconfinement phase transition of $150-170\,$MeV [15, 76].
@@
For the collision energy $\sqrt{s_{NN}}=7.7\,$GeV, we performed calculations for two values of the impact parameters $b=0\,$fm and $b=7.5\,$fm. It can be seen that the values and evolution of the characteristic Unruh temperature do not depend on the centrality of the collision, while the thermodynamic temperature $T$ takes higher values for the central collisions. For the collision energy $\sqrt{s_{NN}}=7.7\,$GeV, the characteristic Unruh temperature averaged over all events and over the entire volume of the fireball is less than the thermodynamic temperature $T$ of the resulting medium at all times after the nuclei touch for both the central ($b=0\,$fm) and off-center ($b=7.5\,$fm) collisions. For $\sqrt{s_{NN}}=11.5\,$GeV at the moment of nuclei touch in a very short time interval ($t-t_{coll}\approx 0.5\,$fm$/c$) the characteristic Unruh temperature $T_{\rm U}$ is greater than the temperature $T$. Note that for $\sqrt{s_{NN}}=11.5\,$GeV, the maximum value of $T_{\rm U}\approx 250\,$MeV exceeds the maximum value of the medium temperature at the moment of maximum nuclei overlap $T\approx 170\,$MeV and corresponds to the acceleration $a\approx 1.6\,$GeV. For a collision energy of $\sqrt{s_{NN}}=4.5\,$GeV in the resulting medium, the characteristic Unruh temperature is significantly less than the temperature $T$ and after the maximum overlap of nuclei ($t-t_{coll}\approx 5\,$fm$/c$) decreases practically to zero.
@@
In the partonic phase, the temperature $T$ is significantly higher than the Unruh temperature $T_{\rm U}$ (Fig. 2 (second panel)), and the latter is very small. In contrast, in the hadronic phase, as can be seen in Fig. 2 (third panel), at the initial moments of time after the collision, the Unruh temperature is significantly higher than $T$. The Unruh temperature rapidly decreases by the time of maximum overlap, but remains higher than $T$ up to times $t-t_{coll}\approx 5\,$fm$/c$. Thus, states with temperature $T$ below the Unruh temperature are found in the hadronic phase, while the medium containing partons is hot and unaccelerated.

### v2

@@ I Introduction
In recent years the field of research of relativistic heavy ion collisions has undergone rapid development, initiated by the active operation of the Relativistic Heavy Ion Collider (RHIC) and, later, of a related program at the Large Hadron Collider (LHC). The discovery of a new state of nuclear matter  [1, 2, 3, 4, 5], known as the quark-gluon plasma (QGP), has prompted a rethinking of the initial assumptions regarding the behavior of high-temperature quantum chromodynamics (QCD) matter. Specifically, the prevailing expectations of a weakly coupled gas-like state [6, 7, 8, 9] have been challenged, suggesting instead a behavior more akin to a strongly-coupled almost ideal fluid [10, 11, 12, 13, 14].
@@
From the point of view of hydrodynamics, the direct analogue of the magnetic field $B_{\mu}=\frac{1}{2}\varepsilon_{\mu\nu\alpha\beta}u^{\nu}F^{\alpha\beta}$ is vorticity $\omega_{\mu}=\frac{1}{4}\varepsilon_{\mu\nu\alpha\beta}u^{\nu}\Omega^{\alpha\beta}$, and the direct analogue of the electric field $E_{\mu}=F_{\mu\nu}u^{\nu}$ is acceleration $a_{\mu}=-\Omega_{\mu\nu}u^{\nu}$ (here $F_{\mu\nu}=\partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu}$ is the usual electromagnetic field tensor, and $\Omega_{\mu\nu}=\partial_{\mu}u_{\nu}-\partial_{\nu}u_{\mu}$ is the kinematic vorticity tensor, while $u^{\mu}$ is the four-velocity of the medium), being, respectively, the vector ($E_{\mu}$ and $a_{\mu}$) and pseudovector ($B_{\mu}$ and $\omega_{\mu}$) components of corresponding tensors ($F_{\mu\nu}$ and $\Omega_{\mu\nu}$). And while vorticity and electric fields have been studied in details, there are not many works devoted to modeling acceleration (see, for example [71, 72]). Of particular interest is the study of acceleration effects using a first-principle lattice approach, since in this case the well-known sign problem can be avoided; the first steps were recently made in this direction [73].
The present research aims to examine the process of generation and evolution of acceleration in both central and non-central heavy-ion collisions simulated within the PHSD model. The focus is on modeling spatial distributions of acceleration at different time moments of the system evolution, with the intention of evaluating acceleration (and the corresponding Unruh temperature) and comparing it with the temperature from the equation of state (EoS). The existence of extremely high accelerated zones, consisting of hadrons, where the Unruh temperature exceeds that obtained from the EoS, will be demonstrated. Furthermore, it will be shown that the partonic phase is characterized by minimal acceleration, which results in the temperature being above the Unruh temperature  [^1] Let us emphasize that the difference between the temperature and the Unruh temperature is not at all connected with any “violation” of the Unruh effect, as it might naively seem. According to the Unruh effect, the temperature is equal to the Unruh temperature only for the concrete quantum (vacuum) state, as discussed below..
@@ II Calculating acceleration and temperature within PHSD
$\quad\times\delta\big(p_{0}-\sqrt{m_{h}^{2}+\boldsymbol{p\,}^{2}}\big)\delta^{(3)}\big(\boldsymbol{r}-\boldsymbol{r}_{i_{h}}(t)\big),$ (3)
@@
$\quad\times\delta\big(p_{0}-\sqrt{m_{h}^{2}+\boldsymbol{p\,}^{2}}\big)\Phi(\boldsymbol{r},\boldsymbol{r}_{i_{h}}(t)),$ (4)
@@
To determine the local temperature, we use the EoS from [89] and solve $T(n_{B},\varepsilon)=T$, where $\varepsilon$ is taken from (6) and $n_{B}$ is the baryon density, calculated from the baryon current $J^{\mu}_{B}$ as $n_{B}=u_{\mu}J^{\mu}_{B}$ [^3] Note that this equation should not be confused with the condition $J^{\mu}_{B}=n_{B}u_{\mu}$, that is, the current is diffusion-less, which fixes the Eckart frame.. Although this EoS is not the most modern, it has proven itself well and has been well tested, in particular, in calculating global polarization. In this paper, we use the hadronic EoS from [89], which does not contain a phase transition to the parton phase  [^4] It should also be noted that PHSD itself has its own built-in equation of state, so the parton phase is described correctly in the kinetic equations. And the equation of state discussed here does not participate in the equations of motion in any way, and is only needed to estimate the temperature.. At the same time, the phase transition is crucial only for the areas of high density and temperature, observed in the first moments after the collision, where, as will be shown, the Unruh temperature will be less than the temperature from EoS (while we are primarily interested just in the areas where the Unruh temperature will be higher than the EoS temperature). And the most difference in the use of other EoS can be only in the area between the core and the corona. Moreover, generally speaking, the temperatures are not much higher than the critical one, which allows us to limit ourselves to this EoS at least at the first stage.
Details of the algorithm used, the explicit form of the smearing kernel $\Phi$, the spectator separation procedure, the analysis of the resulting medium, and the application to the calculation of the global polarization can be found in a series of papers [57, 90, 91, 92].
@@
The PHSD transport model employs the parallel ensemble method, which involves the simultaneous simulation processing of $N$ collisions. To improve the accuracy of the calculated values, the PHSD model is initialized $M$ times, with the fluidization procedure being performed for each run individually. Consequently, the total number of simulated events is $N\times M$. In this paper, the number of parallel ensembles was set to $N=200$ for the collision energies of $7.7$ and 11.511.5\,GeV, and the number of independent model runs is $M\approx 500$. Thus, the total statistics is $(5-10)\times 10^{4}$ events for each impact parameter and collision energy.
@@ III Results
In this sections, we present a detailed analysis of temperature $T$ (obtained from the EoS), the Unruh temperature $T_{\rm U}$ (calculated as (1) and (7)), and the difference $T^{2}-T_{\rm U}^{2}$ for the medium produced in the central (b=0b=0\,fm) and off-center (b=7.5b=7.5\,fm) heavy-ion collisions.
@@ III.1 Temperature and acceleration in central and off-center collisions: overall analysis
Figure 1: Profiles of temperature $T$ (first column), characteristic Unruh temperature $T_{\rm U}$ (second column) in the $z=0$ plane, calculated at different times after the collision for Au + Au collisions at $\sqrt{s_{NN}}=7.7\,$GeV. The profiles of $T$ and $T_{\rm U}$ shown in (a) are calculated for a central collision (b=0b=0\,fm); in (b) for an off-center collision (b=7.5b=7.5\,fm). The dotted curves show two contours in the $xy$ plane, limiting the medium in energy density to the value $\varepsilon=500\,$MeV/fm3 (inner contour) and $\varepsilon=50\,$MeV/fm3 (outer contour).
@@
The time $t-t_{coll}=0\,$fm/c/c corresponds to the moment of contact of the nuclei. At this moment, a hot medium begins to form in the center of the system, which then expands. The volume occupied by the hot substance changes non-trivially [91], and the temperature of the resulting medium reaches its highest value at the moment of maximum overlap of the nuclei ($t-t_{coll}\approx 2.5\,$fm/c/c). The temperature profiles $T$ in Figure 1 show that the hottest medium in the central cut is formed in the time interval from 2 to 3 fm/c/c after the moment of contact of the nuclei for the collision under consideration. After 3 fm/c/c, the resulting medium cools down due to hadronization and thermalization processes [15, 78, 79, 80]. As can be seen from the temperature profiles $T$, the hot medium in the plane z=0z=0\,fm is located inside the contour that limits the medium in energy density to the value $\varepsilon=50\,$MeV/fm3. In what follows, by a fireball we mean a medium limited by this energy density ($\varepsilon\geq 50\,$MeV/fm3). In the PHSD model, the QGP is defined to exist inside a region limited by the energy density to a value of $\varepsilon=500\,$MeV/fm3 from below (see [86] and ref. therein). In Figure 1 we denoted this region by dashed black and white contour. Note that such an analysis is valid for a short period of time after the collision ($t-t_{coll}=3-4\,$fm/c/c). In the PHSD model, after the maximum overlap of nuclei the fraction of the partonic phase rapidly decreases and becomes insignificant [93]. As can be seen in Figure 1, the temperature $T$ reaches a maximum in the region of QGP phase and is about 200 MeV for the time of maximal overlap in the central cut.
@@
The characteristic Unruh temperature, as can be seen from the z=0z=0\,fm profiles in Fig. 1, has a maximal value at early times after the nuclei touch $t-t_{coll}=0.5\,$fm/c/c. Then the value of the characteristic Unruh temperature rapidly decreases with the evolution and expansion of the resulting medium. The most accelerated medium in the central cut is formed mainly outside the fireball, where the medium consists of hadrons. In the QGP region, $T_{\rm U}$ has a very small value at the initial time after the collision and practically disappears by the time of maximum overlap. Thus, in the central cut the cold regions outside the fireball, consisting of hadrons, are highly accelerated. The received pictures exhibit core-corona structures.
@@
The qualitative picture of the evolution of the temperature $T$ and the characteristic Unruh temperature $T_{\rm U}$ remains the same for both central collisions (b=0b=0\,fm, Fig. 1(a)) and off-center ones (b=7.5b=7.5\,fm, Fig. 1(b)). The main difference is that for the non-central collisions the medium volume decreases and the acceleration is slightly higher than for the central ones.
@@ III.2 Comparison: partonic and hadronic phases
To better understand the physical picture presented in Figure 1 at a qualitative level, we calculated a time evolution of the temperatures $T$ and $T_{\rm U}$ separately for partonic and hadronic phases, and averaged over the entire volume of the fireball. Such plots for Au + Au collisions at energies $\sqrt{s_{NN}}=4.5-11.5\,$GeV are presented in Figure 2.
@@
Figure 2: Evolution of the temperature $T$ and the characteristic Unruh temperature $T_{\rm U}$, calculated in PHSD for Au + Au collisions at energies $\sqrt{s_{NN}}=4.5,7.7$ and 11.511.5\,GeV. For the collision energy $\sqrt{s_{NN}}=7.7\,$GeV, two impact parameters are considered: b=0b=0\,fm, which corresponds to a central collision, and b=7.5b=7.5\,fm, which corresponds to an off-center collision of nuclei. The results are presented for a full fireball $\varepsilon>50\,$MeV/fm3 (first panel); for partonic phase only (second panel) and for hadronic phase only (third panel).
@@
Let us first discuss the results obtained for the collision energies $\sqrt{s_{NN}}=7.7$ and 11.511.5\,GeV for the fireball (Fig. 2 (first panel)). After the moment of contact of the nuclei, when the fireball is just beginning to form, the characteristic Unruh temperature has a maximum value. Then the value of $T_{\rm U}$ decreases rapidly, which is associated with the rapid expansion of the fireball. The temperature $T$ of the fireball has a maximum value at the moment of maximum overlap of the nuclei, which corresponds to $t-t_{coll}\approx 2.5\,$fm/c/c. Then the medium begins to cool, which is associated with the processes of hadronization and thermalization. Note that for the collision energies $\sqrt{s_{NN}}=7.7$ and 11.511.5\,GeV, the obtained maximum value of the temperature $T$ of the resulting medium at the moment of maximum overlap is $T\approx 160-170\,$MeV, which is close to the temperature of the confinement-deconfinement phase transition of 150170150-170\,MeV [16, 78].
@@
For the collision energy $\sqrt{s_{NN}}=7.7\,$GeV, we performed calculations for two values of the impact parameters b=0b=0\,fm and b=7.5b=7.5\,fm. It can be seen that the values and evolution of the characteristic Unruh temperature do not depend on the centrality of the collision, while the thermodynamic temperature $T$ takes higher values for the central collisions. For the collision energy $\sqrt{s_{NN}}=7.7\,$GeV, the characteristic Unruh temperature averaged over all events and over the entire volume of the fireball is less than the thermodynamic temperature $T$ of the resulting medium at all times after the nuclei touch for both the central (b=0b=0\,fm) and off-center (b=7.5b=7.5\,fm) collisions. For $\sqrt{s_{NN}}=11.5\,$GeV at the moment of nuclei touch in a very short time interval ($t-t_{coll}\approx 0.5\,$fm/c/c) the characteristic Unruh temperature $T_{\rm U}$ is greater than the temperature $T$. Note that for $\sqrt{s_{NN}}=11.5\,$GeV, the maximum value of $T_{\rm U}\approx 250\,$MeV exceeds the maximum value of the medium temperature at the moment of maximum nuclei overlap $T\approx 170\,$MeV and corresponds to the acceleration $a\approx 1.6\,$GeV. For a collision energy of $\sqrt{s_{NN}}=4.5\,$GeV in the resulting medium, the characteristic Unruh temperature is significantly less than the temperature $T$ and after the maximum overlap of nuclei ($t-t_{coll}\approx 5\,$fm/c/c) decreases practically to zero.
@@
In the partonic phase, the temperature $T$ is significantly higher than the Unruh temperature $T_{\rm U}$ (Fig. 2 (second panel)), and the latter is very small. In contrast, in the hadronic phase, as can be seen in Fig. 2 (third panel), at the initial moments of time after the collision, the Unruh temperature is significantly higher than $T$. The Unruh temperature rapidly decreases by the time of maximum overlap, but remains higher than $T$ up to times $t-t_{coll}\approx 5\,$fm/c/c. Thus, states with temperature $T$ below the Unruh temperature are found in the hadronic phase, while the medium containing partons is hot and unaccelerated.